Suppose that the pressure on a solid is increased quasistatically and isothermally from P to Pr. Show that the change in volume in an infinitesimal step in this process is given bywhere B is the isothermal bulk modulus, and hence that, if variations of V and B with temperature may be neglected, the total work done on the solid by an external agent is
Answer #1

PV=constant

PdV+VdP=0

B=\frac {dP}{\frac {dV}{V}} (Definition of the bulk modulus)

dV=-\left (\frac {V}{B} \right )dP

W_{on}=-PdV=\frac {V}{B}P^2dP

W_{on}=\frac {V}{B}\int_{P_i}^{P_f}P^2dP

W_{on}=\left (\frac {V}{2B} \right )(P_f^2-P_i^2)